The total length of two ships is 1759 feet. The difference in there lengths is 19 feet.A Find the length of each ship. B If a football field has a length of 100yards, determine the length of the longer ship in terms of number of football fields
Found 2 solutions by jorel1380, checkley79:
Answer by jorel1380(3719)

x+x+19=1759
2x=1740
x=870
x+19=889
889ft/300ft=2.96 football fields..

Answer by checkley79(3341)

X+(X+19)=1,759
2X=1,759-19
2X=1,740
X=1,740/2
X=870 FT. FOR THE SHORTER SHIP.
870+19=889 FT. FOR THE LONGER SHIP.
889/100=8.89 FOOTBALL FIELDS.